I could use your help. There are 24 spooks in the haunted house. The number of scarecrows is equal to the number of witches plus 4. The number of witches is equal to the number of ghosts minus 8. how many witches and ghosts and scarecrows are there?
Thanks

Answer by AnlytcPhil(968)

>>...I could use your help. There are 24 spooks in the haunted house. The
number of scarecrows is equal to the number of witches plus 4. The number of
witches is equal to the number of ghosts minus 8. how many witches and ghosts
and scarecrows are there? Thanks
Let x = no. of witches
Let y = no. of ghosts
Let z = no. of scarecrows
>>...There are 24 spooks in the haunted house...<<
Translation: x + y + z = 24
>>...The number of scarecrows is equal to the number of witches plus 4...<<
Translation: z = x + 4
>>...The number of witches is equal to the number of ghosts minus 8...<<
Translation: x = y - 8
So you have this system of three equations and three unknowns:
Can you solve it? If not post again.
x + y + z = 24
z = x + 4
x = y - 8
Answer x=4 witches, y=12 ghosts, z=8 scarecrows
Edwin